OpenStudy (anonymous):

sqrt5^(8n)=125^(n+5)

OpenStudy (anonymous):

Solve fo n

OpenStudy (anonymous):

\[\sqrt{5}^{8n}=125^{n+5}\]

OpenStudy (anonymous):

Thats the problem

OpenStudy (turingtest):

convert\[\sqrt5=5^{1/2}\]and\[125=5^3\]and use the exponent law\[(x^a)^b=x^{ab}\]after that, both sides will have the same base and you can take the logarithm

OpenStudy (anonymous):

Thanks